OpenStudy (anonymous):

the perimeter of a rectangular athletic field is 104m and the length is 16m more than the width. find the length and the width

OpenStudy (anonymous):

p = 2l+2w = 104 l= 16+w 104 = 2(16+w) +2w solve for W